OpenStudy (anonymous):

Find the equilibrium point for the given demand and supply functions. D(p) = (p - 67)2 S(p) = p2 + 12p + 3467 Equilibrium price: pE = Equilibrium quantity: xE =

OpenStudy (anonymous):

Where supply meets demand exist equilibria: $$D(p)=S(p)\\(p-67)^2=p^2+12p+3467\\p^2-134p+4489=p^2+12p+3467\\146p=1022\\p=7$$Our equilibrium price is \(7\), and therefore our quantity is \(S(7)=D(7)=(-60)^2=3600\)